Samples of a cast aluminum part are classified on the basis of surface finish (in microinches) and edge finish. The results of 107 parts are summarized as follows: edge finish excellent good surface excellent 85 finish aood 14 Let A denote the event that a sample has excellent surface finish, and let B denote the event that a sample has excellent edge finish. If a part is selected at random, determine the following probabilities. Round your answers to three decimal places (e.g. 98.765) (a) P(A)10.79 (b) P(B)0.79 c) PA 1- 0.07 (d) P(Ana) (e) P(AUB)= (f) P(A, UB)
Answer #1

a)P(A)=88/107=0.822

b)P(B)=99/107=0.925

c)P(A')=1-0.822=0.178

d)P(A n B)=85/107=0.794

e)P(A u B) =1-5/107=0.953

f)P(A' U B)=1-3/107=0.972